Operations Report – February 7th

SOL: #3
Name of person filing report: Rajvi Patel
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit Rover used: Not used
Opportunity rover used: EVA #3 was postponed due to the windy conditions.
Curiosity rover used: EVA #3 was postponed due to the windy conditions.
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 23.2 gallons
Water (static tank): 416.21 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes
Summary of Internet: Nominal
Summary of suits and radios: Not used
Summary of GreenHab operations: put notes here
WATER USE: 7 gallons
Heater: ON
Supplemental light: 22:00 – 02:00
Harvest: 117g of radish, 23g of Coriander + Basil
Summary of ScienceDome operations: Crew Biologist transferred brought equipment and consumables to the ScienceDome for their project. 140ml of 95% Ethyl Alcohol was used with 60ml of DI water to make 200ml of 70% ethanol for surface sterilization. The fume cabinet was cleaned to make two 5.5ml 1:10 dilutions of brought Flour, SP, and LB media. One of these was kept as a negative control while the other was inoculated with 0.1 grams of gypsum extracted from a larger gypsum crystal from the corner of Brahe Highway and Cow Dung Road from a crew 291 EVA. These were then kept in Oven 1, not turned on, to keep the samples at the internal temperature of 28º C.
Equipment used: Ethanol spray bottle, tweezers, weighing scale (new batteries put in), fume cabinet, cleaning tissues, Oven 1 (not turned on), 100-1000ul pipette, and own pipette tips.
Crew Artist and Journalist used the Science Dome for drawing and painting rock samples for their projects.
Dual split: Heat ON
Summary of RAM operations: The crew Engineer used it to get the tools wire cutter and heavy-duty zip locks to secure locations on the tarp using zip ties. Tools used were wire cutter and heavy-duty zip ties. Used measuring tape to measure the water level. Borrowed another measuring tape for the Crew Biologist’s project in Science Dome.
Now secured
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: put notes here
-Most of the locations were secured using heavy-duty zip ties. However, there are some locations (with torn holes for zip ties to go through) that are required to be secured. (Picture 1 and 2). Advice?
-The photo about securing the bottom of the overlapping tarp section mentioned in the last report is also attached below (highlighted in yellow), along with the type of secured arrangement I would like to propose (highlighted in blue). Advice?

Copyright © The Mars Society. All rights reserved. | Main Site